# Best Arm Identification for Prompt Learning under a Limited Budget

### v2

@@ Appendix A Details of the Adopted BAI-FB Designs
2:  Initialization: set $\hat{\mu}(p)\leftarrow$ for all $p\in{\mathcal{P}}$; set the active prompt set ${\mathcal{A}}\leftarrow{\mathcal{P}}$
@@
2:  Initialization: set $n(p)\leftarrow 0$, $\hat{\mu}(p)\leftarrow$ for all $p\in{\mathcal{P}}$; set the active prompt set ${\mathcal{A}}\leftarrow{\mathcal{P}}$
@@
9:     Compute $p^{\prime}\leftarrow\operatorname*{arg\,min}_{p\in{\mathcal{A}}}\hat{\mu}(p)$ and $\delta_{\tau}\leftarrow\frac{\sum_{p\in{\mathcal{A}}\backslash\{p^{\prime}\}}\hat{\mu}(p)}{|{\mathcal{A}}|-1}-\hat{\mu}(p^{\prime})$
@@
2:  Initialization: set $\hat{\mu}(p)\leftarrow$ for all $p\in{\mathcal{P}}$; set the active prompt set ${\mathcal{A}}\leftarrow{\mathcal{P}}$
@@
Appendix B Full Experimental Details
@@
In this section, we present full details of our experiments. Prior to delving deeper into the discussion, as we use chat-based LLMs, initial system instructions are needed. We adopt the officially recommended system instructions in the experiments, as shown in Fig LABEL:fig:sys_instruction.
B.1 The Adopted Score Functions: Metrics for Evaluation
@@
  • f1-score: Applied to tasks with complex targets like long sentences (e.g., “informal to formal”, “negation”, “sentence similarity”), this metric (defined in Definition B.1), evaluates the overlap between the LLM response and the label.
@@
B.2 Experiments with Fixed Prompt Pools and APE
The prompt generation process to obtain the fixed prompt pools largely follows the one in APE (Zhou et al.,, 2022), i.e., demonstrating LLMs with examples. In particular, in the generation of each prompt, we sample 10 examples from the training set to demonstrate LLMs with two types of generation templates: ‘forward’ and ‘backward’, which are illustrated in Fig. LABEL:fig:generation_temp. The same setups are also adopted in the end-to-end experiments with APE in Sec. 5.2. A side observation is that we find that in general, GPT-3.5 can handle both templates, resulting in reasonable prompts. However, Llama2-7b exhibits difficulties in generating useful prompts from the ‘backward’ template, possibly due to its more simplified structure.
B.3 Experiments with APO
@@
B.4 Implementions of TRIPLE-CLST and TRIPLE-GSE
@@
Appendix C Additional Experimental Results
@@
In this section, additional experimental results are provided to supplement observations in the main paper.
C.1 Additional Details of Clustering
@@
In Sec. 4, clustering is adopted to enhance the efficiency of prompt selection in TRIPLE-CLST. To supplement the two examples provided in Fig. LABEL:fig:cluster, Tables 3 and 4 provide the full details of the prompts to be clustered and showcase prompts that have been grouped into the same clusters. These examples reveal that clustered prompts often share thematic elements or keywords, such as the use of ‘choice/choose’ and ‘movie’ in the movie selection task, Cluster 0, or a focus on ‘rhyming words’ in the rhymes task, Cluster 3. In addition, when the prompt pools are small, we observe that the clustering process struggles to create balanced and meaningful clusters. Thus, the results of TRIPLE-CLS with small prompts pools with $|{\mathcal{P}}|=10$ are not reported. Addressing this challenge could involve developing methods to enforce balanced clustering or innovatively reallocating the budget for more efficient phase I cluster identification. These considerations open avenues for future research.
C.2 Impacts of Expanding Prompt Pools
One additional interesting phenomenon observed is that the expansion of the prompt pool can transform the nature of tasks, oscillating between easy and hard. Specifically, as depicted in Fig. LABEL:fig:change_pool (top), in most tasks, increasing the number of prompts often shifts tasks from easy to hard, i.e., with one clearly better prompt in the pool with size $10$ while having several comparably good prompts in the pool with size $30$. Conversely, as shown in Fig. LABEL:fig:change_pool (bottom), in certain tasks, expanding the prompt pool actually highlights the good prompts.
C.3 Complete Evaluations on $47$ Tasks
@@
In the main paper, we provide experimental results on $12$ representative tasks from the overall $47$ available tasks. In the following, the complete results are discussed.
Instruction-Induction. Results on the $24$ available tasks in the Instruction-Induction dataset (Honovich et al.,, 2022) are illustrated in Fig. 2 (small prompt pools $|{\mathcal{P}}|=10$ with gpt-3.5), 3 (large prompt pools $|{\mathcal{P}}|=30$ with gpt-3.5), and 4 (large prompt pools $|{\mathcal{P}}|=30$ with llam2-7b). It can be observed that TRIPLE methods improve Uniform on most tasks while achieving matching performance on the remaining ones.
BigBench-ii. Results on the $23$ available tasks in the BigBench-ii dataset (Suzgun et al.,, 2022) are reported in Figs. 5 (small prompt pools $|{\mathcal{P}}|=10$ with gpt-3.5), 6 (large prompt pools $|{\mathcal{P}}|=30$ with gpt-3.5), and 7 (large prompt pools $|{\mathcal{P}}|=30$ with Llam2-7b). It can still be observed that TRIPLE methods consistently surpass or at least match the performance of baselines.
@@
Cluster
Prompts
0
The instruction was to select the movie title that appeared the most frequently among the given choices.
@@
3
The instruction was to determine the correct output based on a given input and its corresponding choices.
@@
4
RSelect the correct output film title from the given list of input films.
@@
Cluster
Prompts
0
The instruction was to identify any homophones in the given inputs.
@@
1
The instruction was to find the nearest rhyming word for each given word.
@@
2 The instruction was to identify any words that are still the same after removing the letters ”in”, ”re”, ”pro”, ”ex”, and ”anti” from the beginning or middle of the word.
@@
3
Find the rhyming word for each input word.
@@
4
Replace the word ’phone’ with a similar word.
Identify the words that rhyme with ”phone”.
Identify the words that rhyme with ”phone”.
@@
The instruction was to list the words that rhyme with ”phone”.
5
Provide the correct spelling for the given words.
@@
Replace the letter ”o” with the letter ”a” in each word.
@@
Figure 2: Complete results with small candidate pools $|{\mathcal{P}}|=10$ and gpt-3.5 on the Instruction-Induction dataset under a budget of $50$.
Figure 3: Complete results with large candidate pools $|{\mathcal{P}}|=30$ and gpt-3.5 on the Instruction-Induction dataset under a budget of $150$.
Figure 4: Complete results with large candidate pools $|{\mathcal{P}}|=30$ and llama-7b on the Instruction-Induction dataset under a budget of $150$.
Figure 5: Complete results with small candidate pools $|{\mathcal{P}}|=10$ and gpt-3.5 on the BigBench-ii dataset under a budget of $50$.
Figure 6: Complete results with large candidate pools $|{\mathcal{P}}|=10$ and gpt-3.5 on the BigBench-ii dataset under a budget of $150$.
Figure 7: Complete results with large candidate pools $|{\mathcal{P}}|=30$ and llama-7b on the BigBench-ii dataset under a budget of $150$.

### v3

@@ Appendix C Details of TRIPLE Designs
2:  Initialization: set $\hat{\mu}(p)\leftarrow 0$ for all $p\in{\mathcal{P}}$; set the active prompt set ${\mathcal{A}}\leftarrow{\mathcal{P}}$
@@
2:  Initialization: set $n(p)\leftarrow 0$, $\hat{\mu}(p)\leftarrow 0$ for all $p\in{\mathcal{P}}$; set the active prompt set ${\mathcal{A}}\leftarrow{\mathcal{P}}$
@@
9:     Compute $p^{\prime}\leftarrow\operatorname*{arg\,min}_{p\in{\mathcal{A}}}\hat{\mu}(p)$ and $\delta_{\tau}\leftarrow\min_{p\in{\mathcal{A}}\backslash\{p^{\prime}\}}\{\hat{\mu}(p)-\hat{\mu}(p^{\prime})\}$
@@
2:  Initialization: set $\hat{\mu}(p)\leftarrow 0$ for all $p\in{\mathcal{P}}$; set the active prompt set ${\mathcal{A}}\leftarrow{\mathcal{P}}$
@@
Appendix D Details of TRIPLE’s Extensions to Example Selection
In this section, we provide additional discussions on the extension to example selection mentioned in Sec. 6. It is first noted that the properties of good combinations of examples for few-shot prompts is a complicated topic and an active research problem (Lu et al.,, 2021; Liu et al., 2021a, ; Min et al.,, 2022; Xu and Zhang,, 2024), which still lacks conclusive answers. The proposed designs are based on heuristics that are well recognized and widely evidenced. With deeper understanding of few-shot prompt developed in later research, the perspective provided by TRIPLE and these designs would still be beneficial to guide corresponding modifications and extensions.
CSAR. First, we incorporate the intuitive heuristic that if one example leads to better performance as a one-shot prompt, it contributes positively to the overall few-shot performance (Rubin et al.,, 2022; Li et al.,, 2023). Based on this heuristic, we adapt the CSAR design (Chen et al.,, 2014; Bubeck et al.,, 2013) to perform example selection, which can identify $M$ prompts from ${\mathcal{G}}$ with the highest individual performances, i.e., $\mu(g_{m})$. The details are provided in Alg. 5.
SAR. It is also noticed in previous studies that selecting a diverse set of examples is vital in achieving good few-shot performances (Su et al.,, 2022; Levy et al.,, 2023). Leveraging this heuristic, we propose to first divide the example set ${\mathcal{G}}$ into $M$ clusters, denoted as $\{{\mathcal{G}}^{1},\cdots,{\mathcal{G}}^{M}\}$, based on the embeddings of the examples. Then, for each cluster ${\mathcal{G}}^{m}$, we find one example $g_{m}$ in it with the highest one-shot performance $\mu(g_{m})$. To perform such selection process efficiently, the SAR design (Bubeck et al.,, 2013) is leveraged. In this way, the diversity and quality of the selected examples are both guaranteed. The details are provided in Alg. 6.
Algorithm 5 TRIPLE-CSAR
1:  Input: the set of available examples ${\mathcal{G}}$, the size of the combination $M$, budget $N$
2:  Initialization: set $\hat{\mu}(g)\leftarrow 0$ for all $g\in{\mathcal{G}}$; set the active example set ${\mathcal{A}}\leftarrow{\mathcal{G}}$; $\tilde{T}_{0}\leftarrow 0$; ${\mathcal{G}}_{\text{acc}}\leftarrow\emptyset$; ${\mathcal{G}}_{\text{rej}}\leftarrow\emptyset$; $\tilde{\log}(|{\mathcal{G}}|)\leftarrow\sum_{i\in[|{\mathcal{G}}|]}1/i$
3:  for phase $p=1,\cdots,|{\mathcal{G}}|$ do
4:     $\tilde{T}_{p}\leftarrow\lceil(N-|{\mathcal{G}}|)/(\tilde{\log}(n)(|{\mathcal{G}}|-p+1))\rceil$
5:     Interact with the targeted LLM using each example $g\in{\mathcal{A}}$ as a one-shot prompt for $\tilde{T}_{p}-\tilde{T}_{p-1}$ times
6:     Update the sample means $\{\hat{\mu}(g):g\in{\mathcal{A}}\}$ using the collected samples
7:     Obtain order $\sigma$ such that $\hat{\mu}(g_{\sigma(1)})\geq\hat{\mu}(g_{\sigma(2)})\geq\cdots\geq\hat{\mu}(g_{\sigma(|{\mathcal{A}}|)})$
8:     Compute gaps
$\Delta_{\sigma(r)}\leftarrow\begin{cases}\hat{\mu}(g_{\sigma(r)})-\hat{\mu}(g_{\sigma(M-|{\mathcal{G}}_{\text{acc}}|+1)})&\text{if $r\leq M-|{\mathcal{G}}_{\text{acc}}|$}\\ \hat{\mu}(g_{\sigma(M-|{\mathcal{G}}_{\text{acc}}|)})-\hat{\mu}(g_{\sigma(r)})&\text{if $r\geq M-|{\mathcal{G}}_{\text{acc}}|+1$},\end{cases}\forall r\in[|{\mathcal{A}}|]$
9:     Compute $g^{\prime}\leftarrow\operatorname*{arg\,max}_{g\in{\mathcal{A}}}\Delta_{g}$
10:     Update ${\mathcal{G}}_{\text{acc}}\leftarrow{\mathcal{G}}_{\text{acc}}\cup\{g^{\prime}\}$ if $\hat{\mu}(g^{\prime})\geq\hat{\mu}(g_{\sigma(M-|{\mathcal{G}}_{\text{acc}}|+1)})$; ${\mathcal{G}}_{\text{rej}}\leftarrow{\mathcal{G}}_{\text{acc}}\cup\{g^{\prime}\}$ otherwise
11:     Update ${\mathcal{A}}\leftarrow{\mathcal{G}}/({\mathcal{G}}_{\text{acc}}\cup{\mathcal{G}}_{\text{rej}})$
12:  end for
13:  Output: the set ${\mathcal{G}}_{\text{acc}}$
Algorithm 6 TRIPLE-SAR
1:  Input: the set of available examples ${\mathcal{G}}$ and their embedding ${\mathcal{E}}$, the size of the combination $M$, budget $N$
2:  Cluster ${\mathcal{G}}$ into clusters $\{{\mathcal{G}}^{1},\cdots,{\mathcal{G}}^{M}\}$ based on embeddings ${\mathcal{E}}$ (e.g., via $k$-means)
3:  Initialization: set $\hat{\mu}(g)\leftarrow 0$ for all $g\in{\mathcal{G}}$; set the active example set for cluster $m$ as ${\mathcal{A}}^{m}\leftarrow{\mathcal{G}}^{m}$; set the overall active example set as ${\mathcal{A}}\leftarrow{\mathcal{G}}$; set the active cluster ${\mathcal{M}}\leftarrow[M]$; $\tilde{T}_{0}\leftarrow 0$; $\tilde{\log}(|{\mathcal{G}}|)\leftarrow\sum_{i\in[|{\mathcal{G}}|]}1/i$
4:  for phase $p=1,\cdots,|{\mathcal{G}}|$ do
5:     $\tilde{T}_{p}\leftarrow\lceil(N-|{\mathcal{G}}|)/(\tilde{\log}(n)(|{\mathcal{G}}|-p+1))\rceil$
6:     Interact with the targeted LLM using each example $g\in{\mathcal{A}}$ as a one-shot prompt for $\tilde{T}_{p}-\tilde{T}_{p-1}$ times
7:     Update the sample means $\{\hat{\mu}(g):g\in{\mathcal{A}}\}$ using the collected samples
8:     if $\exists m\in{\mathcal{M}}$ such that $|{\mathcal{A}}^{m}|=1$ then
9:        Update /${\mathcal{M}}\leftarrow{\mathcal{M}}/\{m\}$
10:        Update $g^{*}_{m}\leftarrow$ the remaining example in ${\mathcal{A}}^{m}$
11:     else
12:        $\forall m\in{\mathcal{M}}$, compute $\Delta_{m}\leftarrow\max_{g_{m}\in{\mathcal{A}}^{m}}\{\max_{\bar{g}_{m}\in{\mathcal{A}}^{m}}\hat{\mu}(\bar{g}_{m})-\hat{\mu}(g_{m})\}$ and $g_{m}^{\prime}\leftarrow\operatorname*{arg\,max}_{g_{m}\in{\mathcal{A}}^{m}}\{\max_{\bar{g}_{m}\in{\mathcal{A}}^{m}}\hat{\mu}(\bar{g}_{m})-\hat{\mu}(g_{m})\}$
13:        Compute $m^{\prime}\leftarrow\operatorname*{arg\,max}_{m\in{\mathcal{M}}}\Delta_{m}$
14:        Update ${\mathcal{A}}^{m^{\prime}}\leftarrow{\mathcal{A}}^{m^{\prime}}/\{g^{\prime}_{m^{\prime}}\}$
15:     end if
16:  end for
17:  Output: the set $\{g^{*}_{1},\cdots,g^{*}_{M}\}$
Appendix E Full Experimental Details
In this section, we include full details of our experiments, while the complete codes are also uploaded in the supplementary materials.
E.1 LLM Models and System Instructions
Before further details, we first list the LLM models that we adopted for experiments:
@@ E.1 LLM Models and System Instructions
As we use chat-based LLMs, initial system instructions are needed, where the officially recommended system instructions are adopted in experiments, as shown in Fig LABEL:fig:sys_instruction.
E.2 Score Functions
@@
  • f1-score: Applied to tasks with complex targets like long sentences (e.g., “informal to formal”, “negation”, “sentence similarity”), this metric (defined in Definition E.1), evaluates the overlap between the LLM response and the label.
@@
E.3 Experiments with Fixed Prompt Pools and APE
The prompt generation process to obtain the fixed prompt pools largely follows the one in APE Zhou et al., (2022), i.e., demonstrating LLMs with examples. In particular, in the generation of each prompt, we sample 10 examples from the training set to demonstrate LLMs with two types of generation templates: ‘forward’ and ‘backward’, which are illustrated in Fig. LABEL:fig:generation_temp. The same setups are also adopted in the end-to-end experiments with APE in Sec. 5.2.
A side observation is that we find that in general, GPT-3.5 can handle both templates, resulting in reasonable prompts. However, LLMs with fewer parameter numbers, like Llama2-7b, Gemma-7b, or Mistral-7b-v0.2 we use, exhibit difficulties in generating useful prompts from the ‘backward’ template, possibly due to its more simplified structure.
E.4 Experiments with APO
@@
E.5 Implementions of TRIPLE-CLST and TRIPLE-GSE
@@
E.6 Experiments of Example Selection
In the results reported in Table 4, the task is to select $4$ examples from a set of overall $50$ candidate examples within $100$ interactions with the targeted LLM. For tasks with a training dataset larger than 50 examples, $50$ examples are first sampled to construct the candidate set, which is used consistently across experiments. There are also a few tasks with a training dataset smaller than $50$ examples, which are thus entirely used as the candidate set. The same prompt template, as illustrated in Fig. LABEL:fig:few_shot, is used for all tasks to maintain consistency.
Further details regarding the adopted baselines are provided in the following.
  • Random. To validate the benefits of interactions with the targeted LLM during the selection, one commonly adopted baseline is to randomly select the required number of examples from the candidate pool.
  • Uniform. Similar to the uniform baseline adopted in Sec. 5, the overall budget can be uniformly divided to evaluate the one-shot performance of each prompt. Then, the examples with the highest estimated one-shot performances are selected.
Also, for TRIPLE-SAR, the same process of obtaining embeddings as described in Appendix E.5 with also $k$-means as the algorithm to perform clustering.
E.7 Computing Resources and Costs
We use a workstation with two Nvidia-A6000 Ada GPUs for all experiments using white-box LLMs (i.e., Llama2, Mistral, and Gemma). To reproduce our result, any GPU with over 30 GB of memory should be sufficient. With our equipment, each interaction with the white-box LLMs typically takes around $1.3-2.0$ seconds. For experiments using GPT-3.5, the whole execution is light regarding local computational resources, while accesses to the OpenAI API is needed to perform learning. Under our network condition, one API call typically takes around $1$ second.
Appendix F Additional Experimental Results
@@
Additional experimental results are provided to supplement observations in the main paper.
F.1 Additional Details of Clustering
@@
In Sec. 4, clustering is adopted to enhance the efficiency of prompt selection in TRIPLE-CLST. Fig. LABEL:fig:cluster is provided to demonstrate the effectiveness of clustering, i.e., prompts with similar performances are often clustered together. To supplement the two examples provided in Fig. LABEL:fig:cluster, Tables 5 and 6 provide the full details of the prompts to be clustered and showcase prompts that have been grouped into the same clusters. These examples reveal that clustered prompts often share thematic elements or keywords, such as the use of ‘choice/choose’ and ‘movie’ in the movie selection task, Cluster 0, or a focus on ‘rhyming words’ in the rhymes task, Cluster 3.
F.2 Selection of Budgets
To further guide practical implementation, we additionally investigate how to select a reasonable budget. In particular, we focus on the efficiency of various prompt selection algorithms in identifying a “good” prompt – either the optimal prompt in the pool or achieving $95\%$ or better of the optimal prompt’s performance. Fig. 5 illustrates that initial increases in budgets significantly improve the probability of identifying a good prompt, but this benefit tapers off with further budget expansion. This finding suggests that starting with a modest budget and incrementally increasing it is the more effective approach, stopping when additional investment no longer translates into significant returns.
Figure 5: Probability for different algorithms to select a good prompt under different budgets (right), collected with GPT-3.5 and averaged over $5$ runs.
F.3 Performances on Gemma and Mistral
For the experiments on the selected tasks with $|{\mathcal{P}}|=30$ prompts and budget $N=150$, additional results with Gemma and Mistral are reported in Fig. LABEL:fig:gemma_30 and LABEL:fig:mistral_30. The superiority of TRIPLE can still be observed, demonstrating its flexibility over different LLMs.
F.4 Complete Evaluations on $47$ Tasks
@@
In the main paper, we provide experimental results of $12$ representative tasks from the overall $47$ available tasks in Sec. 5. In the following, the complete results are discussed.
  • $|{\mathcal{P}}|=30$, $N=150$: results on the $24$ available tasks in the Instruction-Induction dataset (Honovich et al.,, 2022) are illustrated in Fig. LABEL:fig:all_induction_30 (GPT-3.5), and LABEL:fig:all_induction_llama_30 (Llama2);
  • $|{\mathcal{P}}|=30$, $N=150$: results on the $23$ available tasks in the BigBench-ii dataset (Suzgun et al.,, 2022) are illustrated in Fig. LABEL:fig:all_bigbench_30 (GPT-3.5), and LABEL:fig:all_bigbench_llama_30 (Llama2).
  • $|{\mathcal{P}}|=150$, $N=100$: results on the $24$ available tasks in the Instruction-Induction dataset (Honovich et al.,, 2022) are illustrated in Fig. LABEL:fig:all_induction_150 (GPT-3.5), and LABEL:fig:all_induction_llama_150 (Llama2);
  • $|{\mathcal{P}}|=150$, $N=100$: results on the $23$ available tasks in the BigBench-ii dataset (Suzgun et al.,, 2022) are illustrated in Fig. LABEL:fig:all_bigbench_150 (GPT-3.5), and LABEL:fig:all_bigbench_llama_150 (Llama2).
Also, the complete performances of example selection for few-shot prompts discussed in Sec. 6 are presented in Fig. 7 (Instruction-Induction) and 8 (BigBench-ii).
@@
Cluster Prompts
0 The instruction was to select the movie title that appeared the most frequently among the given choices.
@@
3 The instruction was to determine the correct output based on a given input and its corresponding choices.
@@
4 RSelect the correct output film title from the given list of input films.
@@
Cluster Prompts
0 The instruction was to identify any homophones in the given inputs.
@@
1 The instruction was to find the nearest rhyming word for each given word.
@@
2 The instruction was to identify any words that are still the same after removing the letters "in", "re", "pro", "ex", and "anti" from the beginning or middle of the word.
@@
3 Find the rhyming word for each input word.
@@
4 Replace the word ’phone’ with a similar word.
Identify the words that rhyme with "phone".
Identify the words that rhyme with "phone".
@@
The instruction was to list the words that rhyme with "phone".
5 Provide the correct spelling for the given words.
@@
Replace the letter "o" with the letter "a" in each word.
@@
Figure 7: Complete few-shot results on the Instruction-Induction dataset using GPT-3.5 with $|{\mathcal{G}}|=50$ examples, budget $N=100$, and prompt length $M=4$.
Figure 8: Complete few-shot results on the Instruction-Induction dataset using GPT-3.5 with $|{\mathcal{G}}|=50$ examples, budget $N=100$, and prompt length $M=4$.